OpenStudy (geekfromthefutur):

PLZ Help I will fan and give a medal Which exponential function goes through the points (1, 16) and (4, 128)? f(x) = 4(4)x f(x) = 8(2)x f(x) = 8(2)-x f(x) = 4(4)-x

OpenStudy (anonymous):

@geekfromthefutur are the "x"s exponents?

OpenStudy (geekfromthefutur):

Yes

OpenStudy (anonymous):

Okay, then this is quite a simple problem let me walk you through it. The basic equation for exponential graphs are y=ab^x By plugging in both points into that equation we can solve for a and b respectively. So we plug in (1,16) into the equation 16 = ab^1 We also plug in (4, 128) into the equation 128 = ab^4 If you multiply 16 by 8 we get 128, so multiply 8 to both sides. 128 = 8ab^1 128 = ab^4 128 = 128 8ab^1 = ab^4 8ab = ab^4 Divide both sides by a 8b = b^4 Divide both sides by b 8 = b^3 Cube root both sides 2 = b Choose one of the equations we made and plug b in. 16 = ab^1 16 = ab 16 = 2a a = 8 SO your exponential equation will be y = 8(2)^x Second answer choice. :)

OpenStudy (anonymous):

Hopefully that made sense, haha

OpenStudy (geekfromthefutur):

Dang i was wondering what was takin so long bro

OpenStudy (anonymous):

Well, I wanted to make sure you understood properly. :p

OpenStudy (geekfromthefutur):

Yea i do now